given pointP(-3,5),S(1,-2),R(5,6)&Q(5,1)find the gradient of PQ & RS
Answer by stanbon(75887)

given pointP(-3,5),S(1,-2),R(5,6)&Q(5,1)find the gradient of PQ & RS
-----
PQ gradient = slope = (1-5)/(5--3) = -4/8 = -1/2
----
RS gradient = slope = (6--2)/(5-1) = 8/4 = 2
------------------
Cheers,
Stan H.
--------------